. In a population of 200 individuals, 72 are homozygous recessive for the character of eye color (cc). One hundred individuals from this population die from a fatal disease. Thirty-six of the survivors are homozygous recessive. In the original population, the frequency of the dominant allele is ???
Answer #1

N = population size = 200

q^2 = 72/200 = 0.36;

q = 0.6;

p = 1-q = 1-0.6 = 0.4 (= frequency of dominant allele in the original population).

p^2 = 0.16.

Now, 100 individuals are dead. So, population size becomes 100.

q^2=36/100 =0.36; q=0.6; p = 0.4